OpenStudy (j3nny0rtiz):

please help me You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly?

Directrix (directrix):

Binomial Probability Problem with n = 5 and P(correct) = 1/4

OpenStudy (maddielb):

@j3nny0rtiz I think its 45/512

Directrix (directrix):

>>I think its 45/512 How did you get that? @MaddieLB

Directrix (directrix):

P(x = 3) = C(5,3) * (1/4)^3 * (3/4)^2 What does this equal? C(5,3) * (1/4)^3 * (3/4)^2 = ?

OpenStudy (maddielb):

I dont even know lol, I just threw something together. You'd probably know better than me

OpenStudy (j3nny0rtiz):

I'm confuse